# Dequantified Diffusion-Schrödinger Bridge for Density Ratio Estimation

### v3

@@ B.1 Special Cases of DI
In the case of TRE, the interpolation strategy, as detailed in Eq. (2), represents a specific case of $\mathbf{I}$, characterized by $\alpha_{t}=\sqrt{1-\eta_{t}^{2}},\beta_{t}=\eta_{t}$, with $t$ taking discrete values $10,1/M,2/M,\ldots,$ 1. For DRE-$\infty$, the coefficients are defined as $\alpha_{t}=\exp\{-0.25(\beta_{\max}-\beta_{\min})t^{2}-0.5\beta_{\min}t\}$ and $\beta_{t}=\sqrt{1-\alpha_{t}^{2}}$ for the MNIST dataset and $\alpha_{t}=1-t$ and $\beta_{t}=t$ for other datasets. The corresponding stochastic process for the former one aligns with the solution to variance preserving (VP) SDEs (Song et al., 2021; Li et al., 2024b; Xin et al., 2024; Zhou et al., 2025b).
@@ C.1 Comparison of the Trajectory Sets for Interpolation Strategies
In this section, we provide a detailed comparison of interpolation strategies, specifically deterministic interpolant (DI), diffusion bridge interpolant (DBI), dequantified diffusion bridge interpolant (DDBI), and dequantified Schrödinger bridge interpolant (DSBI). Their intermediate samples and corresponding distributions are visualized in Figure 1.
(a) DI: DI constrains intermediate samples to fixed linear paths between $q_{0}(\mathbf{x})$ and $q_{1}(\mathbf{x})$, resulting in narrow bands across the trajectory space (Figure 1(a)). While dense along the paths, DI severely limits support and fails to explore alternative trajectories, making it inflexible and unsuitable for diverse distributions.
(b) DBI: DBI introduces stochasticity through Brownian Bridge noise, expanding support and enabling broader trajectory exploration (Figure 1(b)). Compared to DI, DBI provides greater coverage and variability while retaining tractability, reducing the rigidity of interpolation paths (Zhao et al., 2023).
(c) DDBI: Extending DBI, DDBI modulates the noise with deterministic interpolation weights and diffusion components. This results in more dispersed trajectories (Figure 1(c)) and a larger coverage of the intermediate distributions, balancing controlled stochasticity with enhanced flexibility (Zhou et al., 2025a).
(d) DSBI: DSBI offers full stochastic control over noise and leverages entropy-regularized optimal transport, resulting in widely dispersed trajectories and efficient utilization of the trajectory space (Figure 1(d)). By minimizing transition loss, DSBI achieves the largest support set and highest diversity among the methods, producing rich intermediate distributions.
@@ C.2 Joint Score Matching
In this section, we integrate the time score $s_{\boldsymbol{\theta}}^{t}\in\mathbb{R}$ and data score $\mathbf{s}_{\boldsymbol{\theta}}^{\mathbf{x}}\in\mathbb{R}^{d}$ to formulate the joint score $\mathbf{s}_{\boldsymbol{\theta}}^{t,\mathbf{x}}:[s_{\boldsymbol{\theta}}^{t},\mathbf{s}_{\boldsymbol{\theta}}^{\mathbf{x}}]\in\mathbb{R}^{d+1}.$ This joint score is incorporated into the training objective defined in Equation 23, resulting in a joint score matching objective (Choi et al., 2022):
@@ C.3 Mutual Information Estimation
We train the score model using the joint score matching loss (details in Section C.2). The batch size is set to 512 for $d=\{40,80,160\}$ and 256 for $d=320$, with iteration steps of $\{40k,100k,400k,500k\}$, respectively. DRE-$\infty$ serves as the baseline method. Results, shown in Figure 3, demonstrate that $\text{D}^{3}\text{RE}$, especially DSBI, produces MI estimates significantly closer and faster to the ground truth compared to the baseline, highlighting its superiority in accurately capturing mutual dependencies between variables.
@@ C.4 Density Estimation
2-D Synthetic Datasets. In this section, we present density estimation results on eight synthetic datasets for different methods. From left to right, the epochs are 0, 2000, 4000, 6000, 8000, 10000, 12000, 14000, 16000, 18000 and 20000. Corresponding results are shown in Figures 9, 10, 11, 12, 13, 14, 15 and 16. D3RE (including DDBI and DSBI) achieved the best performance on all datasets and was able to learn the best results with fewer epochs.
@@ C.5 Ablation Study on $\gamma^{2}$
[…]
We also present density estimation results on eight synthetic datasets for varing values of $\gamma^{2}$. From left to right, the epochs are 0, 2000, 4000, 6000, 8000, 10000, 12000, 14000, 16000, 18000 and 20000. Corresponding results are shown in Figures 17, 18, 19, 20, 21, 22, 23 and 24.
@@ C.6 Ablation Study on OTR
Figure 3 compares the MI estimation performance of DDBI and DSBI across different dimensions ($d=80,120$). DDBI uses diffusion bridges and Gaussian dequantization, while DSBI adds OTR to achieve better alignment. In panel (a), both DDBI and DSBI outperform the baseline methods. DDBI shows stable performance and converges close to the ground truth. However, DSBI, with OTR, achieves faster convergence and higher accuracy, staying closer to the ground truth throughout training. In panel (b), the impact of OTR becomes more evident. Although DDBI still outperforms the baseline methods, its convergence is slower, and its accuracy is lower compared to DSBI. By leveraging OTR, DSBI demonstrates superior MI estimation performance across all training epochs and remains closer to the ground truth in high-dimensional settings ($d=120$). OTR significantly improves the alignment of intermediate distributions and enhances model performance. When combined with diffusion bridges and Gaussian dequantization, as in DSBI, OTR achieves its full potential. It allows the model to estimate complex distributions more accurately.
Number of Function Evaluations. We analyze the impact of OTR on NFE, noting that DI and DDBI do not utilize OTR. Our observations show that applying OTR significantly reduces NFE. Figure 8 compares NFE across four methods in DRE, highlighting substantial variations in computational efficiency. The first approach exhibits the highest NFE, indicating reliance on iterative procedures requiring repeated function evaluations. The second approach achieves a moderate reduction in NFE, likely by minimizing redundant evaluations through minimized transport costs.

### v5

@@ B.1 Special Cases of DI
In the case of TRE, the interpolation strategy, as detailed in Eq. (2), represents a specific case of $\mathbf{I}$, characterized by $\alpha_{t}=\sqrt{1-\eta_{t}^{2}},\beta_{t}=\eta_{t}$, with $t$ taking discrete values $0,1/M,2/M,\ldots,1$. For DRE-$\infty$, the coefficients are defined as $\alpha_{t}=\exp\{-0.25(\beta_{\max}-\beta_{\min})t^{2}-0.5\beta_{\min}t\}$ and $\beta_{t}=\sqrt{1-\alpha_{t}^{2}}$ for the MNIST dataset and $\alpha_{t}=1-t$ and $\beta_{t}=t$ for other datasets. The corresponding stochastic process for the former one aligns with the solution to variance preserving (VP) SDEs (Song et al., 2021; Li et al., 2024b; Xin et al., 2024; Zhou et al., 2025b).
@@ C.1 Comparison of the Trajectory Sets for Interpolation Strategies
In this section, we provide a detailed comparison of interpolation strategies, specifically deterministic interpolant (DI), diffusion bridge interpolant (DBI), dequantified diffusion bridge interpolant (DDBI), and dequantified Schrödinger bridge interpolant (DSBI). Their intermediate samples and corresponding distributions are visualized in Fig. 1.
(a) DI: DI constrains intermediate samples to fixed linear paths between $q_{0}(\mathbf{x})$ and $q_{1}(\mathbf{x})$, resulting in narrow bands across the trajectory space (Fig. 1(a)). While dense along the paths, DI severely limits support and fails to explore alternative trajectories, making it inflexible and unsuitable for diverse distributions.
(b) DBI: DBI introduces stochasticity through Brownian Bridge noise, expanding support and enabling broader trajectory exploration (Fig. 1(b)). Compared to DI, DBI provides greater coverage and variability while retaining tractability, reducing the rigidity of interpolation paths (Zhao et al., 2023).
(c) DDBI: Extending DBI, DDBI modulates the noise with deterministic interpolation weights and diffusion components. This results in more dispersed trajectories (Fig. 1(c)) and a larger coverage of the intermediate distributions, balancing controlled stochasticity with enhanced flexibility (Zhou et al., 2025a).
(d) DSBI: DSBI offers full stochastic control over noise and leverages entropy-regularized optimal transport, resulting in widely dispersed trajectories and efficient utilization of the trajectory space (Fig. 1(d)). By minimizing transition loss, DSBI achieves the largest support set and highest diversity among the methods, producing rich intermediate distributions.
@@ C.2 Joint Score Matching
In this section, we integrate the time score $s_{\boldsymbol{\theta}}^{t}\in\mathbb{R}$ and data score $\mathbf{s}_{\boldsymbol{\theta}}^{\mathbf{x}}\in\mathbb{R}^{d}$ to formulate the joint score $\mathbf{s}_{\boldsymbol{\theta}}^{t,\mathbf{x}}:[s_{\boldsymbol{\theta}}^{t},\mathbf{s}_{\boldsymbol{\theta}}^{\mathbf{x}}]\in\mathbb{R}^{d+1}.$ This joint score is incorporated into the training objective defined in Eq. 24, resulting in a joint score matching objective (Choi et al., 2022):
@@ C.3 Mutual Information Estimation
We train the score model using the joint score matching loss (details in Sec. C.2). The batch size is set to 512 for $d=\{40,80,160\}$ and 256 for $d=320$, with iteration steps of $\{40k,100k,400k,500k\}$, respectively. DRE-$\infty$ serves as the baseline method. Results, shown in Fig. 3, demonstrate that $\text{D}^{3}\text{RE}$, especially DSBI, produces MI estimates significantly closer and faster to the ground truth compared to the baseline, highlighting its superiority in accurately capturing mutual dependencies between variables.
@@ C.4 Density Estimation
2-D Synthetic Datasets. In this section, we present density estimation results on eight synthetic datasets for different methods. From left to right, the epochs are 0, 2000, 4000, 6000, 8000, 10000, 12000, 14000, 16000, 18000 and 20000. Corresponding results are shown in Figs. 9, 10, 11, 12, 13, 14, 15 and 16. D3RE (including DDBI and DSBI) achieved the best performance on all datasets and was able to learn the best results with fewer epochs.
@@ C.5 Ablation Study on $\gamma^{2}$
Mutual Information Estimation. The ablation study on varying $\gamma^{2}$ values (Fig. 5) reveals distinct convergence behaviors in MI estimation across epochs. For all dimensions ($d=\{40,80,120\}$), smaller $\gamma^{2}$ values ($\leq 0.01$) lead to faster initial convergence toward the DRE-$\infty$ baseline, particularly in lower dimensions ($d=40$). However, excessively small $\gamma^{2}=0.001$ introduces instability in later epochs, causing slight deviations from the baseline. In contrast, larger $\gamma^{2}$ values ($\geq 0.1$) show slower initial convergence but stabilize over longer training periods, especially in higher dimensions ($d=120$). Notably, $\gamma^{2}=0.1$ strikes a balance between convergence speed and stability, consistently aligning with the baseline across all dimensions. These findings suggest that the optimal $\gamma^{2}$ selection is influenced by both the dimensionality and training duration, with moderate regularization ($\gamma^{2}=0.01$$0.1$) providing robust MI estimation performance.
Density Estimation. The ablation study on $\gamma^{2}$ for density estimation (Fig. 4) reveals systematic trade-offs in performance across regularization strengths. For small $\gamma^{2}=0.001$, the model achieves rapid initial alignment with the ground truth distribution (first row) but exhibits overfitting artifacts in later epochs, manifesting as irregular density peaks and deviations from the smooth ground truth structure. Intermediate values ($\gamma^{2}=0.01$$0.1$) demonstrate balanced behavior: $\gamma^{2}=0.01$ preserves finer details while maintaining stability, and $\gamma^{2}=0.1$ produces smoother approximations with minimal divergence from the true distribution. Larger $\gamma^{2}$ values ($\geq 0.5$) induce excessive regularization, leading to oversmoothed estimates that fail to capture critical modes of the 2-D data, particularly in high-density regions. Notably, $\gamma^{2}=0.1$ achieves the closest visual and structural resemblance to the ground truth, suggesting its suitability for low-dimensional tasks requiring both fidelity and robustness. These results underscore the necessity of tuning $\gamma^{2}$ to mitigate under-regularization artifacts while preserving distributional complexity.
We also present density estimation results on eight synthetic datasets for varing values of $\gamma^{2}$. From left to right, the epochs are 0, 2000, 4000, 6000, 8000, 10000, 12000, 14000, 16000, 18000 and 20000. Corresponding results are shown in Figs. 17, 18, 19, 20, 21, 22, 23 and 24.
@@ C.6 Ablation Study on OTR
Fig. 3 compares the MI estimation performance of DDBI and DSBI across different dimensions ($d=80,120$). DDBI uses diffusion bridges and Gaussian dequantization, while DSBI adds OTR to achieve better alignment. In panel (a), both DDBI and DSBI outperform the baseline methods. DDBI shows stable performance and converges close to the ground truth. However, DSBI, with OTR, achieves faster convergence and higher accuracy, staying closer to the ground truth throughout training. In panel (b), the impact of OTR becomes more evident. Although DDBI still outperforms the baseline methods, its convergence is slower, and its accuracy is lower compared to DSBI. By leveraging OTR, DSBI demonstrates superior MI estimation performance across all training epochs and remains closer to the ground truth in high-dimensional settings ($d=120$). OTR significantly improves the alignment of intermediate distributions and enhances model performance. When combined with diffusion bridges and Gaussian dequantization, as in DSBI, OTR achieves its full potential. It allows the model to estimate complex distributions more accurately.
Number of Function Evaluations. We analyze the impact of OTR on NFE, noting that DI and DDBI do not utilize OTR. Our observations show that applying OTR significantly reduces NFE. Fig. 8 compares NFE across four methods in DRE, highlighting substantial variations in computational efficiency. The first approach exhibits the highest NFE, indicating reliance on iterative procedures requiring repeated function evaluations. The second approach achieves a moderate reduction in NFE, likely by minimizing redundant evaluations through minimized transport costs.